# MARS: Unleashing the Power of Variance Reduction for Training Large Models

### v2

@@ 5 Experiments
In this section, we evaluate the performances of two instantiations of our algorithm, MARS-AdamW and MARS-Lion [^2] For the sake of training efficiency, we use MARS-approx for the experiments as the default configuration, except for Appendices D.2 and D.4. Discussion of the difference in performance between MARS-exact and MARS-approx is in Appendix D.2., in comparison with AdamW (Loshchilov and Hutter, 2019), the predominant algorithm for training large language models, Lion (Chen et al., 2023) and Muon (Jordan et al., 2024) on GPT-2 model series. More experiment results and abaltion study, including the computer vision experiments, the effect of different learning rate schedulers, as well as sensitivity to $\gamma$, are postponed to Section D.
@@ 5.1 Experimental Setup
All our experiments are done based on the nanoGPT (Karpathy, 2022) implementation of the GPT-2 (Radford et al., 2019) architecture, and on the OpenWebText (Gokaslan et al., 2019) dataset. The training and validation sets contain approximately $9$ billion and $4.4$ million tokens, respectively, all preprocessed using the GPT-2 tokenizer. We conduct experiments on three scales of GPT-2 models: small (125M parameters), medium (355M parameters), and large (770M parameters). Per the nanoGPT configurations, we disabled biases, applied GeLU activations, and set the Dropout rate (Srivastava et al., 2014) to $0.0$. We utilized 16 NVIDIA A100 GPUs for training the small models. For the medium and large models, training was conducted on 32 NVIDIA A100 GPUs and 32 NVIDIA H100 GPUs, respectively. Other hyper-parameters of training are listed in Appendix E.
@@ 5.2 Results
In Figure 1 and Figures 67 (in the Appendix), we demonstrate the training and validation losses as a function of training tokens and wall-clock time for various model sizes [^3] The training loss curves are smoothed using Exponential Moving Average.. Across the small, medium, and large GPT-2 models, MARS consistently surpasses both the AdamW and Muon baselines in training and validation losses. The performance gap becomes more pronounced with increasing model size. Notably, MARS exhibits both rapid initial decay and sustained superiority throughout the training process. Further, we explore the performance of additional learning rate choices in Appendix D. Notably, the best validation losses of MARS-AdamW and MARS-Lion achieved in our GPT-2 large experiments are 2.511 and 2.534. For comparison, the best validation losses are 2.568, 2.565 and 2.606 for AdamW, Lion and Muon. These results demonstrate that our reported performance is highly competitive with state-of-the-art optimizers.
In Figure 1(c), as well as Figures 6(c) and 7(c) in the Appendix, we compare the wall-clock time of different algorithms. We observe that MARS-AdamW and MARS-Lion have a slightly higher per-iteration cost compared to AdamW but is much faster than Muon. Additionally, they consistently demonstrate lower validation losses than both AdamW, Muon and Lion within equivalent training durations.
We also evaluate 0-shot and 5-shot performances of our optimizer on common benchmarks including ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020) and MMLU (Hendrycks et al., 2021), with the lm-evaluation-harness codebase (Gao et al., 2024). We only list the 5-shot performances for large models in Table 1, and leave other results in the Appendix D.1. The models pre-trained with MARS-AdamW and MARS-Lion outperform those pre-trained with AdamW, Muon and Lion optimizers, validating an enhanced downstream performance within the same number of pre-training steps.
5.3 Different Learning Rate Scheduler
5.3.1 Constant LR
To ensure a fair comparison by eliminating the influence of learning rate changes during training and to explore the potential for continuous training with MARS, we conduct supplementary experiments on GPT-2 small, medium, and large models using a constant learning rate for both AdamW and MARS-AdamW-approx. For each group of experiments, we compared between 2 different maximum learning rates. The training and validation curves are displayed in Figures 2 and 3. The results also indicate that our algorithm has superior performances over AdamW optimizer under a fair comparison of constant learning rates.
5.3.2 WSD Scheduler
Cosine learning rate scheduler is utilized in most of our experiences. Recently Hu et al. (2024) introduced a novel learning rate scheduler called Warmup-Stable-Decay (WSD) scheduler, which composed of 3 stages, including learning rate linear-warmup stage, constant learning rate stage, as well as learning rate decay stage. To test the flexibility and ability of continuous training of MARS trained with respect to learning rate schedulers, we implement experiments on GPT-2 small and medium models with AdamW and MARS-AdamW-approx scheduled with WSD for 10k, 20k, 50k and 100k steps. The training and validation loss curves are shown in Figures 4 and 5. The curves indicate that MARS has a better potential for continuous training and exhibits explicit edge over baseline algorithm.
@@ 6 Conclusion
Acknowledgments
We would like to thank Yiming Dong, Zhuoqing Song, Yuan Cao, Volkan Cevher, Francesco Orabona, Laurence Aitchison, Yida Yin, Zhuang Liu and Zhiyuan Li for their valuable comments and feedback on our work.
@@ Appendix A Related Work
RProp (Riedmiller and Braun, 1993) is probably one of the earliest adaptive gradient methods by dynamically adjusting the learning rate. AdaGrad (Duchi et al., 2011; McMahan and Streeter, 2010) adjusts the learning rate based on the geometry of the training data observed during earlier iterations. To tackle with the issue of diminishing gradient in AdaGrad (Carlson et al., 2015a, b), Tieleman (2012) introduced RMSProp by incorporating the idea of exponential moving average. A significant advancement came with Adam (Kingma and Ba, 2015), which integrated RMSProp with Nesterov’s momentum (Nesterov, 1983, 2013) achieving superior performance and becoming a prevalent optimizer in deep neural network training. Later, Loshchilov and Hutter (2019) proposed to decouple weight decay from gradient calculations in Adam and introduced AdamW, an optimization algorithm having become the predominant optimization algorithm in contemporary deep learning applications. To fix the convergence issue of Adam, Reddi et al. (2019b) introduced the AMSGrad optimizer, which maintains a running maximum of past second-order momentum terms to achieve non-increasing step sizes. Subsequently, Chen et al. (2018a) unified AMSGrad and SGD within the Padam framework by introducing a partial adaptive parameter to control the degree of adaptiveness. Notably, AdamW and its variations have been widely used in the training of popular large language models, including OPT (Zhang et al., 2022a), Llama 3 (Dubey et al., 2024), and DeepSeek-V3 (Liu et al., 2024).
Variance Reduction Methods.
@@
SAG(Roux et al., 2012) and SDCA(Shalev-Shwartz and Zhang, 2013) were among the first attempts to apply variance reduction techniques to accelerate the convergence of SGD. Subsequently, simpler algorithms like SVRG(Johnson and Zhang, 2013) and SAGA(Defazio et al., 2014) were introduced, achieving the same improved convergence rates. SARAH (Nguyen et al., 2017a) further simplified these approaches by employing biased recursive gradient estimation, which reduces storage requirements while achieving the complexity bounds for convex optimization problems. For non-convex optimization, besides SVRG (Allen-Zhu and Yuan, 2016; Reddi et al., 2016) and SARAH (Nguyen et al., 2017b), SPIDER (Fang et al., 2018) integrates Normalized Gradient Descent (Nesterov, 2013; Hazan et al., 2015) with recursive estimation of gradients, while SNVRG (Zhou et al., 2020) introduces multiple reference points for semi-stochastic gradient calculation for improved variance reduction and convergence rate. SpiderBoost (Wang et al., 2019) refines SPIDER by enabling the use of a significantly larger constant step size while preserving the same near-optimal oracle complexity. Subsequently, STORM (Cutkosky and Orabona, 2019) was proposed to further simplifies the SPIDER and SNVRG algorithms through the use of stochastic recursive momentum. This was later improved into a parameter-free variant, namely STORM+(Levy et al., 2021).
Variance Reduction for Adaptive Gradient Methods.
@@
Few works have explored the application of variance reduction techniques to adaptive gradient methods. To the best of our knowledge, the only exceptions are Adam+, SuperAdam and AdaSPIDER. Adam+ (Liu et al., 2020) attempts to reduce the variance of first-order moment into Adam by estimating the gradient only at extrapolated points. SuperAdam (Huang et al., 2021) integrates variance reduction with AdamW to achieve improved convergence rates. And AdaSPIDER (Kavis et al., 2022) introduced adaptive step size in SPIDER algorithm. However, these variance-reduced adaptive gradient methods have primarily been validated on basic computer vision tasks, such as MNIST (Schölkopf and Smola, 2002) and CIFAR-10 (Krizhevsky et al., 2009), and simple natural language modeling tasks, like SWB-300 (Saon et al., 2017), using straightforward architectures such as LeNet (LeCun et al., 1998), ResNet-32 (He et al., 2016), 2-layer LSTMs (Graves and Graves, 2012), and 2-layer Transformers (Vaswani, 2017). As a result, a significant gap remains in the successful application of variance reduction techniques to adaptive gradient methods, particularly in the rapidly evolving domain of large language models.
Appendix B Proof of Theorems

### v3

@@ 4 Experiments
In this section, we evaluate the performances of two instantiations of our algorithm, MARS-AdamW and MARS-Lion [^2] For the sake of training efficiency, we use MARS-approx for the experiments as the default configuration, except for Appendices E.2 and E.4. Discussion of the difference in performance between MARS-exact and MARS-approx is in Appendix E.2., in comparison with AdamW (Loshchilov & Hutter, 2019), the predominant algorithm for training large language models, Lion (Chen et al., 2023) and Muon (Jordan et al., 2024) on GPT-2 model series. More experiment results and abaltion study, including the computer vision experiments, the effect of different learning rate schedulers, as well as sensitivity to $\gamma$ and batch size, are postponed to Section E.
@@ 4.1 Experimental Setup
All our experiments are done based on the nanoGPT (Karpathy, 2022) implementation of the GPT-2 (Radford et al., 2019) architecture, and on the OpenWebText (Gokaslan et al., 2019) dataset. The training and validation sets contain approximately $9$ billion and $4.4$ million tokens, respectively, all preprocessed using the GPT-2 tokenizer. We conduct experiments on three scales of GPT-2 models: small (125M parameters), medium (355M parameters), and large (770M parameters). Per the nanoGPT configurations, we disabled biases, applied GeLU activations, and set the Dropout rate (Srivastava et al., 2014) to $0.0$. We utilized 16 NVIDIA A100 GPUs for training the small models. For the medium and large models, training was conducted on 32 NVIDIA A100 GPUs and 32 NVIDIA H100 GPUs, respectively. Other hyper-parameters of training are listed in Appendix F.
@@ 4.2 Results
In Figure 1 and Figures 23 (in the Appendix), we demonstrate the training and validation losses as a function of training tokens and wall-clock time for various model sizes [^3] The training loss curves are smoothed using Exponential Moving Average.. Across the small, medium, and large GPT-2 models, MARS consistently surpasses both the AdamW and Muon baselines in training and validation losses. The performance gap becomes more pronounced with increasing model size. Notably, MARS exhibits both rapid initial decay and sustained superiority throughout the training process. Further, we explore the performance of additional learning rate choices in Appendix E. Notably, the best validation losses of MARS-AdamW and MARS-Lion achieved in our GPT-2 large experiments are 2.511 and 2.534. For comparison, the best validation losses are 2.568, 2.565 and 2.606 for AdamW, Lion and Muon. These results demonstrate that our reported performance is highly competitive with state-of-the-art optimizers.
In Figure 1(c), as well as Figures 2(c) and 3(c) in the Appendix, we compare the wall-clock time of different algorithms. We observe that MARS-AdamW and MARS-Lion have a slightly higher per-iteration cost compared to AdamW but is much faster than Muon. Additionally, they consistently demonstrate lower validation losses than both AdamW, Muon and Lion within equivalent training durations.
We also evaluate 0-shot and 5-shot performances of our optimizer on common benchmarks including ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020) and MMLU (Hendrycks et al., 2021), with the lm-evaluation-harness codebase (Gao et al., 2024). We only list the 5-shot performances for large models in Table 1, and leave other results in the Appendix E.1. The models pre-trained with MARS-AdamW and MARS-Lion outperform those pre-trained with AdamW, Muon and Lion optimizers, validating an enhanced downstream performance within the same number of pre-training steps.
@@ 5 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of optimization theory in deep learning. We believe that our work contributes meaningfully to the field, specifically on advancing the efficiency in the pre-training stage of deep learning models, especially Large Language Models. By involvement of variance reduction in adaptive learning methods, our method can greatly lower the cost for pre-training language models on limited training corpora and more resource-constrained devices as well as in broader settings, opening new avenues for their application in various downstream tasks. Improvement in efficiency typically correlates with reduced energy consumption, potentially decreasing the environmental footprint of LLM pre-training. This advancement underscores the potential of optimization method development in deep learning field in both technological and societal contexts.
@@ Appendix A Related Work
Adaptive Gradient Methods. RProp (Riedmiller & Braun, 1993) is probably one of the earliest adaptive gradient methods by dynamically adjusting the learning rate. AdaGrad (Duchi et al., 2011; McMahan & Streeter, 2010) adjusts the learning rate based on the geometry of the training data observed during earlier iterations. To tackle with the issue of diminishing gradient in AdaGrad (Carlson et al., 2015a, b), Tieleman (2012) introduced RMSProp by incorporating the idea of exponential moving average. A significant advancement came with Adam (Kingma & Ba, 2015), which integrated RMSProp with Nesterov’s momentum (Nesterov, 1983, 2013) achieving superior performance and becoming a prevalent optimizer in deep neural network training. Later, Loshchilov & Hutter (2019) proposed to decouple weight decay from gradient calculations in Adam and introduced AdamW, an optimization algorithm having become the predominant optimization algorithm in contemporary deep learning applications. To fix the convergence issue of Adam, Reddi et al. (2019b) introduced the AMSGrad optimizer, which maintains a running maximum of past second-order momentum terms to achieve non-increasing step sizes. Subsequently, Chen et al. (2018) unified AMSGrad and SGD within the Padam framework by introducing a partial adaptive parameter to control the degree of adaptiveness. Notably, AdamW and its variations have been widely used in the training of popular large language models, including OPT (Zhang et al., 2022a), Llama 3 (Dubey et al., 2024), and DeepSeek-V2 (Liu et al., 2024).
@@
Variance Reduction Methods. SAG(Roux et al., 2012) and SDCA(Shalev-Shwartz & Zhang, 2013) were among the first attempts to apply variance reduction techniques to accelerate the convergence of SGD. Subsequently, simpler algorithms like SVRG(Johnson & Zhang, 2013) and SAGA(Defazio et al., 2014) were introduced, achieving the same improved convergence rates. SARAH (Nguyen et al., 2017a) further simplified these approaches by employing biased recursive gradient estimation, which reduces storage requirements while achieving the complexity bounds for convex optimization problems. And some researchers have also attempted to apply preconditioning into variance reduction in the convex setting (Frangella et al., 2024; Derezinski, ). For non-convex optimization, besides SVRG (Allen-Zhu & Yuan, 2016; Reddi et al., 2016) and SARAH (Nguyen et al., 2017b), SPIDER (Fang et al., 2018) integrates Normalized Gradient Descent (Nesterov, 2013; Hazan et al., 2015) with recursive estimation of gradients, while SNVRG (Zhou et al., 2020) introduces multiple reference points for semi-stochastic gradient calculation for improved variance reduction and convergence rate. SpiderBoost (Wang et al., 2019) refines SPIDER by enabling the use of a significantly larger constant step size while preserving the same near-optimal oracle complexity. Subsequently, STORM (Cutkosky & Orabona, 2019) was proposed to further simplifies the SPIDER and SNVRG algorithms through the use of stochastic recursive momentum. This was later improved into a parameter-free variant, namely STORM+(Levy et al., 2021).
@@
Variance Reduction for Adaptive Gradient Methods. Few works have explored the application of variance reduction techniques to adaptive gradient methods. To the best of our knowledge, the only exceptions are Adam+, SuperAdam and AdaSPIDER. Adam+ (Liu et al., 2020) attempts to reduce the variance of first-order moment into Adam by estimating the gradient only at extrapolated points. SuperAdam (Huang et al., 2021) and VRAdam (Li, 2024) integrates variance reduction with AdamW to achieve improved convergence rates. And AdaSPIDER (Kavis et al., 2022) introduced adaptive step size in SPIDER algorithm. However, these variance-reduced adaptive gradient methods have primarily been validated on basic computer vision tasks, such as MNIST (Schölkopf & Smola, 2002) and CIFAR-10 (Krizhevsky et al., 2009), and simple natural language modeling tasks, like SWB-300 (Saon et al., 2017), using straightforward architectures such as LeNet (LeCun et al., 1998), ResNet-32 (He et al., 2016), 2-layer LSTMs (Graves & Graves, 2012), and 2-layer Transformers (Vaswani, 2017). As a result, a significant gap remains in the successful application of variance reduction techniques to adaptive gradient methods, particularly in the rapidly evolving domain of large language models.
Appendix B Theoretical Analysis
B.1 Connection to Nesterov’s Acceleration
Many optimization algorithms exhibit similarities with Nesterov’s acceleration and can be considered adaptations of Nesterov’s acceleration with varying parameterization schedules, as discussed in works such as Defazio et al. (2024) and Xie et al. (2024). In this section, we compare and contrast Nesterov’s momentum and STORM momentum used in our paper. Specifically, Nesterov’s accelerated gradient descent can be equivalently written as (Xie et al., 2024):
$\mathbf{m}_{t}=\beta_{1}\mathbf{m}_{t-1}+\big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\beta_{1}(\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{]},$
while the STORM momentum is
$\mathbf{m}_{t}=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\beta_{1}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}$
The most significant difference lies in the noise handling schemes. In Nesterov’s acceleration, $\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ is subtracted by $\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})$ to determine a direction of improvement. In contrast, STORM variance reduction subtracts $\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})$ from $\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ to cancel out the noise introduced by $\bm{\xi}_{t}$. Furthermore, in our theoretical analysis (Section B.2), we prove that variance-reduced variants of AdamW achieve an improved convergence rate of $\mathcal{O}(T^{-1/3})$. Empirically, we also show that a variance reduced noise schedule performs better than its approximate counterpart.
B.2 Convergence of MARS
Although Kingma & Ba (2015) did convergence analysis for Adam, Reddi et al. (2019a) pointed out that they made some mistakes in the proof, and they also proved that in some special cases, Adam does not converge. However, there are some attempts to prove the convergence of Adam and AdamW in special circumstances (Zhang et al., 2022b; Li et al., 2024; Zhou et al., 2024). Our algorithm, MARS, is also based on AdamW. In addition, it involves the property of variance reduction. We prove that MARS can converge with a better convergence rate with careful selection of hyperparameters.
To help analyze the convergence of the algorithm, we make the following assumptions:
Assumption B.1 (Bounded Variance).
We assume that the variance of gradient estimator is bounded by $\sigma^{2}$. i.e., for any noise $\bm{\xi}$, parameter $\bm{x}$, and $\nabla F(\bm{x})=\mathbb{E}[\nabla f(\bm{x},\bm{\xi})]$, there exists a positive $\sigma$ such that:
$\mathbb{E}\big{[}\|\nabla f(\mathbf{x},\bm{\xi})-\nabla F(\mathbf{x})\|_{2}^{2}\big{]}\leq\sigma^{2}.$ (B.1)
Assumption B.2 ($L$-Smoothness).
We assume that for arbitrary $\bm{\xi}$, $f(\bm{x},\bm{\xi})$ is $L$-smooth:
$\|\nabla f(\mathbf{x},\bm{\xi})-\nabla f(\mathbf{y},\bm{\xi})\|_{2}\leq L\|\mathbf{x}-\mathbf{y}\|_{2},~{}\forall\mathbf{x},\mathbf{y}.$ (B.2)
Assumption B.3 ($H$ Lower Bounded).
We assume that there is a constant $\rho>0$ such that for all $\mathbf{H}_{t},t>0$, $\mathbf{H}_{t}\succ\rho\bm{I}$.
Remark B.4.
Note that this is implicitly satisfied by the instantiations of MARS since we add a small $\epsilon$ to semi-positive definite $\mathbf{H}_{t}$ for computational stability.
We proposed Theorem B.5 for our main Algorithm 1 and Theorem B.6 for MARS-AdamW (Algorithm 2, where an additional weight decay is involved). We note that for theoretical analysis, it is necessary to consider time-varying parameters $\beta_{1,t}$ and $\beta_{2,t}$. However, in practice, these parameters are typically set as constants.
Theorem B.5.
In Algorithm 1, under Assumptions B.1B.2 and B.3, when choosing $\eta_{t}=(s+t)^{-1/3},s\geq 8L^{3}/\rho^{3}$. Suppose $c\geq 32L^{2}\rho^{-2}+1$, $\beta_{1,t+1}=1-c\eta_{t}^{2}$ and $\beta_{2,t+1}=1-\eta_{t}^{6}$, then $\forall T\geq s$, it holds that
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}$ $\leq\Big{(}2\rho G+\frac{\rho c^{2}\sigma^{2}}{4L^{2}}\cdot\log(s+T)\Big{)}\cdot\frac{1}{T^{2/3}}-\frac{\rho^{2}\sum_{t=1}^{T}M_{t+1}}{8L^{2}T^{1/3}},$
$\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}$ $\leq\Big{(}\frac{16G}{3\rho}+\frac{2c^{2}\sigma^{2}}{3L^{2}}\cdot\log(s+T)\Big{)}\cdot\frac{1}{T^{2/3}}-\frac{\sum_{t=1}^{T}M_{t+1}}{6L^{2}T^{1/3}},$
where $G=F(\mathbf{x}_{1})-\min_{\mathbf{x}}F(\mathbf{x})+\frac{\rho s^{1/3}\sigma^{2}}{16L^{2}}$ and $M_{t+1}$ is defined in (C.2).
Theorem B.6.
In Algorithm 2, under Assumptions B.1B.2 and B.3, when choosing $\eta_{t}=(s+t)^{-1/3},s\geq\max(8L^{3}/\rho^{3},64\lambda^{3})$. Suppose $||\mathbf{x}_{t}||_{2}\leq D$, $c\geq 32L^{2}\rho^{-2}+1$, $\beta_{1,t+1}=1-c\eta_{t}^{2}$ and $\beta_{2,t+1}=1-\eta_{t}^{6}$, then $\forall T\geq s$, it holds that
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}$ $\leq\Big{(}2\rho(G+\lambda D^{2}\log(s+T))+\frac{\rho c^{2}\sigma^{2}}{4L^{2}}\cdot\log(s+T)\Big{)}\cdot\frac{1}{T^{2/3}}-\frac{\rho^{2}\sum_{t=1}^{T}M_{t+1}}{16L^{2}T^{1/3}},$
$\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\eta_{t}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}$ $\leq\Big{(}\frac{16(G+\lambda D^{2}\log(s+T))}{\rho}+\frac{2c^{2}\sigma^{2}}{L^{2}}\cdot\log(s+T)\Big{)}\cdot\frac{1}{T^{2/3}}-\frac{\sum_{t=1}^{T}M_{t+1}}{L^{2}T^{1/3}}.$
where $G=F(\mathbf{x}_{1})-\min_{\mathbf{x}}F(\mathbf{x})+\frac{\lambda}{2}D^{2}(1+\epsilon)+\frac{\rho s^{1/3}\sigma^{2}}{16L^{2}}$ and $M_{t+1}$ is defined in (C.2).
The theorems above guarantee the convergence rate of $O(\log(T)/T^{1/3})$. We remark that even though it seems that the involvement of weight decay may results in slower convergence, it performs better in practice, à la Loshchilov & Hutter (2019). Finally, the term $M_{t+1}$ is always greater or equal to 00, and when $\gamma_{t+1}=1$, we would have $M_{t+1}=0$. The dependency of $M_{t+1}$ on $\gamma_{t+1}$ is shown in (C.2) and Lemma C.2. This proves that the convergence is faster if we allow a flexible $\gamma_{t}$ schedule.
Appendix C Proof of Theorems